Consider the reaction of tin and bromine. Sn2+ + Br2 → Sn4+ + 2Br-. The concentration of tin (ii) is 0.050M and the tin (iv) and bromide concentrations is 0.00010M
Q = [ Sn4+][Br-]2 / [Sn2+] and Sn2+ / Sn4+ = +0.150; Br2/Br- = +1.065
Find the voltage of the electrochemical cell using the Nernst equation.
1
Expert's answer
2019-08-22T03:22:10-0400

E=(EBr2/2Br--ESn4+/Sn2+)-(0,059/n)*lg((asn2+)/(a2Br-*aSn4+))=(1,065-0,150)-(0,059/2)*lg(0,050/0,0001030)=0,599V

E=0,599V